OpenStudy (anonymous):

2 times the square root x -5 = 2

OpenStudy (anonymous):

7

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf 2 \sqrt{x-5}=2 }\) divide both sides by 2 \(\Huge\color{blue}{ \sf \sqrt{x-5}=1 }\) square both sides \(\Huge\color{blue}{ \sf (\sqrt{x-5})^2=(1)^2 }\) (simplify...) \(\Huge\color{blue}{ \sf x-5=1 }\) add 5 to both sides \(\Huge\color{blue}{ \sf x-5+5=1 +5}\) \(\Huge\color{blue}{ \sf x=6 }\)

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \frak Welcome~~to }\) \(\Huge\color{orangered}{ \frak OPENSTUDY! }\)

OpenStudy (anonymous):

why would u divide ?

OpenStudy (solomonzelman):

both sides by 2? To eliminate the square root

OpenStudy (solomonzelman):

to isolate not eliminate (sorry)

OpenStudy (solomonzelman):

\(\large\color{blue}{ \it If~~you~~~are~~not~~clear~~Ask! }\)

OpenStudy (anonymous):

so x= 6 so would the solution be not extraneous?

OpenStudy (solomonzelman):

\(\Huge\color{red}{ \sf 2\sqrt{x-5}=2 }\) \(\Huge\color{red}{ \sf 2\sqrt{\color{royalblue} { 6 }-5}=2 }\) \(\Huge\color{red}{ \sf 2\sqrt{1}=2 }\) \(\Huge\color{red}{ \sf 2=2 }\) SO....... YES OR NO ?

OpenStudy (anonymous):

I have to choose one so is it no?

OpenStudy (solomonzelman):

it is yes. because when i plugged it in it worked.

OpenStudy (anonymous):

yeah I got wrong but thanks.

OpenStudy (solomonzelman):

WHAT?! you got it wrong?!?! But that's impossible !

OpenStudy (solomonzelman):

6 is for sure correct, I don't know what they teach in schools, but this I DO know ... agh I am so mad bro ! How could that be !!!!!!!!!!!!!??????

OpenStudy (anonymous):

it probably was the part where it asked whether it was extraneous or not. but, its ok thanks though.